# GNN-DT: A Graph Neural Network Enhanced Decision Transformer for Efficient Optimization in Dynamic Environments

### v2

@@ 1 Introduction
Sequential decision-making problems are critical for efficiently operating a wide array of industries, such as power systems control [1], logistics optimization [2], portfolio management [3], and advanced manufacturing processes [4]. However, many practical problems, such as the electric vehicle (EV) charging optimization [5], are large-scale, have temporal dependencies, and aggregated constraints, often making conventional methods impractical [6]. This is especially observed in dynamic environments, where the optimization landscape continuously evolves, requiring real-time solutions.
@@
Reinforcement learning (RL) [7] has been extensively studied for solving optimization problems due to its ability to manage uncertainty, adapt to dynamic environments, and enhance decision-making through trial-and-error [8, 9]. In complex and large-scale scenarios, RL can provide high-quality solutions in real-time compared to traditional mathematical programming techniques that fail to do so [10]. However, RL approaches face significant challenges, such as sparse reward signals that slow learning and hinder convergence to optimal policies [11]. In addition, RL solutions struggle to generalize when deployed in environments different from the one they were trained in, limiting their applicability in real-world scenarios with constantly changing conditions [12].
@@
Decision Transformers (DT) [13] is an offline RL algorithm that reframes traditional RL problems as generative sequence modeling tasks conditioned on future rewards [14]. By learning from historical data, DTs effectively address the sparse reward issue inherent in online RL, relying on demonstrated successful outcomes instead of extensive trial-and-error exploration. However, the trajectory-stitching mechanism of DT often proves insufficient in dynamic real-world environments, leading to suboptimal policies. Although improved variants such as Q-regularized DT (Q-DT) [15] incorporate additional constraints for greater robustness, they still face significant challenges in generalizing across non-stationary tasks [16]. Consequently, further architectural advances and training strategies are essential to ensure consistent performance in complex environments.
@@
This study introduces GNN-DT [^1] The code can be found at https://github.com/StavrosOrf/DT4EVs. , a novel DT architecture that leverages the permutation-equivariant properties of Graph Neural Networks (GNNs) to handle dynamically changing state-action spaces (i.e., varying numbers of nodes over time) and improve generalization. By generating embeddings that remain consistent under node reordering, GNNs offer a powerful way to capture relational information in complex dynamic environments. Moreover, GNN-DT features a novel residual connection between input and output tokens, ensuring that action outputs are informed by the dynamically learned state embeddings for more robust decision-making. To demonstrate the superior performance of the proposed method, we conduct extensive experiments on the complex multi-objective EV charging optimization problem [17], which encompasses sparse rewards, temporal dependencies, and aggregated constraints. The main contributions are summarized as follows:
@@ 2 Related Work
Advancements in Decision Transformers
@@
Classic DT encounters significant challenges, including limited trajectory stitching capabilities and difficulties in adapting to online environments. To address these issues, several enhancements have been proposed. The Q-DT [15] improves the ability to derive optimal policies from sub-optimal trajectories by relabeling return-to-go values in the training data. Elastic DT [18] enhances classic DT by enabling trajectory stitching during action inference at test time, while Multi-Game DT [19] advances its task generalization capabilities. The Online DT [20, 21] extends DTs to online settings by combining offline pretraining with online fine-tuning, facilitating continuous policy updates in dynamic environments. Additionally, adaptations for offline safe RL incorporate cost tokens alongside rewards [22, 23]. DT has also been effectively applied to real-world domains, such as healthcare [14] and chip design [24], showcasing its practical utility.
RL for EV Smart Charging
@@
RL algorithms offer notable advantages for EV dispatch, including the ability to handle nonlinear models, robustly quantify uncertainty, and deliver faster computation than traditional mathematical programming [25]. Popular methods, such as DDPG [26], SAC [27], and batch RL [28], show promise but often lack formal constraint satisfaction guarantees and struggle to scale with high-dimensional state-action spaces [29, 30]. Safe RL frameworks address these drawbacks by imposing constraints via constrained MDPs, but typically sacrifice performance and scalability [31, 32]. Multiagent RL techniques distribute complexity across multiple agents, e.g., charging points, stations, or aggregators [33], yet still encounter convergence challenges and may underperform in large-scale applications. To the best of our knowledge, no study has used DTs to solve the complex EV charging problem, despite DTs’ potential to handle sparse rewards effectively.
3 Preliminaries
3.1 Offline RL
Offline RL aims to learn a policy $\pi_{\theta}(a\!\mid\!s)$ that maximizes the expected discounted return $\mathbb{E}\bigl{[}\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\bigr{]}$ without additional interactions with the environment [34]. A Markov Decision Process (MDP) is defined by the tuple $(S,A,P,R,\gamma)$, where $S$ is the state space, $A$ the action space, $P$ the transition function, $R$ the reward function, $\gamma\in(0,1]$ the discount factor [7]. In the offline setting, a static dataset $\mathcal{D}=\{(s,a,r)\}$, collected by a (potentially suboptimal) policy, is provided. DTs leverage this dataset by treating RL trajectories as sequences, learning to predict actions that maximize returns based on previously collected experiences. A key component in DTs is the return-to-go (RTG), which for a time step $t$ can be defined as: $G_{t}=\sum_{\tau=t}^{T}\gamma^{\tau-t}\,r_{\tau},$ representing the discounted cumulative reward from $t$ until the terminal time $T$. Offline RL is particularly beneficial when real-time exploration is costly or impractical, while sufficient historical data are available.
3.2 The EV Smart Charging Problem
Working closely with a charge point operator (CPO), it was evident that existing heuristic and mathematical programming charging strategies don’t scale efficiently as EV fleets grow. To address this, we designed the state–action space and objectives around real-world operational constraints and assumptions provided by the CPO. We consider a set of $\mathcal{I}$ charging stations indexed i, all assumed to be controlled by a CPO over a time window $\mathcal{T}$, divided into $T$ non-overlapping intervals. For a given time window, each charging station $i$ operates a set of $\mathcal{J}$ non-overlapping charging sessions, denoted by $\mathcal{J}_{i}=\{j_{1,i},\cdots,j_{J_{i},i}\}$, where $j_{j,i}$ represents the $j^{th}$ charging event at the $i^{th}$ charging station and $J_{i}=|\mathcal{J}_{i}|$ is the total number of charging sessions seen by charging station i in an episode. A charging session is then represented as $j_{j,i}:\{t^{a}_{j,i},t^{d}_{j,i},\bar{p}_{j,i},e^{*}_{j,i}\},~{}\forall j,i$, where $t^{a},t^{d},\bar{p}$ and $e^{*}$ represent the arrival time, departure time, maximum charging power, and the desired battery energy level at the departure time. The primary goal is to minimize the total energy cost given by:
Maintaining the desired battery charge at departure is crucial for EV user satisfaction. We model as:
Eq. (3) defines a sparse reward added at each EV departure based on its departure energy level. Building on these objectives, the overall EV charging problem is formulated as:
$\max_{p^{+},p^{-}}\left(f_{1}(\cdot)-100\cdot f_{2}(\cdot)-10\cdot f_{3}(\cdot)\right)$ (4)
The multi-objective optimization function in Eq.4 integrates Eqs.13 using experimentally determined coefficients based on practical importance. This mixed integer programming (MIP) problem is subject to lower-level operational constraints (e.g., EV battery, power levels) as detailed in AppendixA.1.
EV Charging MDP
The optimal EV charging problem can be framed as an MDP: $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},R)$. At any time step $t$, the state $\boldsymbol{s}_{t}\in\mathcal{S}$ is represented by a dynamic graph $\mathcal{G}_{t}=(\mathcal{N}_{t},\mathcal{E}_{t})$, where $\mathcal{N}_{t}$ is the set of nodes and $\mathcal{E}_{t}$ is the set of edges. The graph is dynamic since the number of nodes in the state and action graph can vary in each step, because of EVs’ arrival and departures. Each node $n\in\mathcal{N}_{t}$ has a feature vector $\boldsymbol{x}_{n}^{t}\in\mathbb{R}^{d}$, capturing node-dependent information such as power limits and prices (see Appendix A.3 for the complete definition). This graph structure [35] efficiently models evolving relationships among EVs, chargers, and the grid infrastructure. The action space $\mathbf{a}_{t}\in\mathcal{A}$ is represented by a dynamic graph $\mathcal{G}^{\mathbf{a}}_{t}=(\mathcal{N}^{\mathbf{a}}_{t},\mathcal{E}^{\mathbf{a}}_{t})$, where nodes $\mathcal{N}^{\mathbf{a}}_{t}$ correspond to the decision variables of the optimization problem (e.g., EVs). Each node $n\in\mathcal{N}^{\mathbf{a}}_{t}$ represents a single action $a_{i,t}\in\mathbf{a}_{t}$, scaled by the corresponding charger’s maximum power limit. For charging, $a_{i,t}\in[0,1]$, and for discharging, $a_{i,t}\in[-1,0)$. The transition function $\mathcal{P}(\boldsymbol{s}_{t+1}\mid\boldsymbol{s}_{t},\mathbf{a}_{t})$ accounts for uncertainties in EV arrivals, departures, energy demands, and grid fluctuations. Finally, the reward function is the same with the objective described in Eq. 4 and is defined as $R(\boldsymbol{s}_{t},\mathbf{a}_{t})=f_{1}-100f_{2}-10f_{3}$ for a single timestep ($\mathcal{T}=\{t\}$). The reward is guiding the policy to maximize cost savings, respect operational constraints, and meet EV driver requirements. While Eqs.1 and 2 represent individual EV rewards and aggregated EV penalties, respectively, Eq.3 introduces a sparse reward that activates only when an EV departs, thereby creating complex temporal dependencies. The reward function weights were selected after comprehensive experimentation, ensuring the optimal solution of the MIP remains the same while the RL training is balanced between exploration and penalization.
4 GNN-based Decision Transformer
4.1 Sequence Embeddings
In detail, the State Embedder consists of $L$ consecutive Graph Convolutional Network (GCN) [36] layers, which aggregate information from neighboring nodes as follows:
4.2 Decoding Actions
4.3 Training: Action Masking and Loss Function
The proposed GNN-DT model is trained via supervised learning using an offline trajectory dataset [13], similarly to offline RL. Specifically, the GPT-2 model is initialized with its default pre-trained weights, which are subsequently fine-tuned end-to-end for the EV charging optimization task. In GNN-DT, the learning of infeasible actions, such as charging an unavailable EV, is avoided through action masking. At each time step $t$, a mask vector $\mathbf{m}_{t}$, which has the same dimension as $\mathbf{a}_{t}$, is generated with zeros marking invalid actions and ones marking valid actions. For example, an action is invalid when the $a_{i,t}\neq 0$ and no EV is connected at charger $i$. The mean squared error between the predicted actions $\widehat{\mathbf{a}}_{t}$ and ground-truth actions $\mathbf{a}_{t}$ from expert or offline trajectories is employed as the loss function. For a window of length $K$ ending at time $t$, training loss is defined as:
5 Experiments
Experimental Setup
The dataset generation and the evaluation experiments are conducted using the EV2Gym simulator [17], which leverages real-world data distributions, including EV arrivals, EV specifications, electricity prices, etc. This setup ensures a realistic environment where the state and action spaces accurately reflect real charging stations’ operational complexity. A scenario with 25 chargers is chosen, allowing up to 25 EVs to be connected simultaneously. In this configuration, the action vector has up to 25 variables (one per EV), while the state vector contains around 150 variables describing EV statuses, charger conditions, power transformer constraints, and broader environmental factors. Consequently, the resulting optimization problem is in the moderate-to-large scale range, reflecting the key complexities of real-world EV charging. Each training procedure is repeated 10 times with distinct random seeds to ensure statistically robust findings. All reported rewards represent the average performance over 50 evaluation scenarios, each featuring different configurations (electricity prices, EV behavior, power limits, etc.). Training was carried out on an NVIDIA A10 GPU paired with 11 CPU cores and 80 GB of RAM, using the AdamW optimizer and a LambdaLR scheduler. Baseline RL agents converged in 2–5 hours, while the proposed GNN-DT required up to 24 hours of training. Default hyperparameters were used for all baseline RL methods; Appendix B.1 lists the full set of hyperparameters employed to train the DTs.
Dataset Generation
Offline RL algorithms, including DTs, can learn policies from trajectories without needing online interaction with the environment. Consequently, the quality of the gathered training trajectories has a substantial impact on the learning process. In this work, three distinct strategies were used to generate trajectories: a) Random: Uniformly sampled actions in the range $[-1,1]$ were applied to the simulator. b) Business-as-Usual (BaU): A Round Robin heuristic charging policy commonly employed by CPOs, which sequentially allocates charging power among EVs to balance fairness and efficiency. c) Optimal: Expert solutions derived from solving offline the MIP as described in Section 3.2 for multiple scenarios. Each trajectory consists of 300 state-action-reward-action mask tuples, with each timestep representing a 15-minute interval, resulting in a total of three simulated days. This combination of random, typical, and expert data provides a comprehensive basis for evaluating how GNN-DT learns from diverse offline trajectories.
5.1 Training Performance
Fig. 2 compares the proposed GNN-DT against multiple baselines, including the classic DT [13] and Q-DT [15] , which both rely on flattened state representations due to their inability to directly process graph-structured data. In these baseline methods, empty chargers and unavailable actions are replaced by zeros, so the action vector is always the same size. Several well-known online RL algorithms from the Stable-Baselines-3 [37] framework (SAC, DDPG, TD3, TRPO, PPO, and TQC) and offline RL algorithms from D3RLPY [38] (IQL, CQL, and BC) are also included. The offline RL algorithms (IQL, CQL, BC, DT, Q-DT, and GNN-DT) are trained on three datasets (Optimal, Random, and BaU), each comprising 10.000 trajectories. A red dotted line marks the optimal reward, which represents the experimental maximum achievable reward obtained by solving the deterministic MIP knowing the future (Oracle) defined in Appendix A.1. This oracle reward serves as an upper bound and helps contextualize the relative performance of each method. It is important to note that the EV smart charging problem requires real-time (1-5 min intervals) optimization solutions for large-scale, highly stochastic scenarios, where metaheuristic algorithms, stochastic optimization, and model predictive control methods fail due to computational constraints. By contrast, once trained (over 2–24 hours), RL agents can deliver real-time charging schedules on an ordinary computer in milliseconds.
In Figs. 2.a-c, the DT-based approaches use a context length $K=10$. As expected, the Optimal dataset provides the highest-quality information, enabling GNN-DT to converge rapidly toward near-oracle performance, while classic DT, Q-DT, and the other offline RL algorithms lag far behind, showcasing GNN-DT’s improved sample efficiency. With the Random dataset, the limited quality of data leads all methods to plateau at lower reward values, although GNN-DT still surpasses the other baselines. An intriguing behavior is observed with the BaU dataset, where classic DT, BC, and IQL converge at rewards exceeding that of GNN-DT. In contrast, the online RL algorithms displayed in Fig. 2.d struggle to achieve comparable improvements, suggesting that pure online exploration is insufficient for solving this complex EV charging optimization problem with sparse rewards. For completeness, Appendices B.2 and B.3 contain the training curves for all algorithm-dataset-context length configurations used. In the rest of this section, we omit the online and offline RL baselines, as their performance is substantially inferior to that of DT, Q-DT, and GNN-DT.
Dataset Impact
In Table 1, the maximum episode reward is compared for small, medium, and large datasets (100, 1.000, and 10.000 trajectories), under two different context lengths ($K=2$ and $K=10$). The left side of Table 1 reports the dataset type, the number of trajectories, and the average reward in each dataset. All baselines achieve performance above the Random dataset’s average reward. However, only GNN-DT consistently approaches the Optimal dataset’s performance, reaching as close as $-0.04\times 10^{5}$ compared to the $-0.01\times 10^{5}$ optimal reward. This advantage becomes especially evident at the largest dataset size (10.000 trajectories), highlighting the benefits of the graph-based embedding layer. Overall, GNN-DT outperforms the baselines across all datasets and both context lengths, with the single exception of the BaU dataset at $K=10$. Interestingly, a larger context window does not always translate into higher rewards, potentially due to the problem setting. Similarly, the dataset size appears to have minimal impact on Q-DT, whereas DT and GNN-DT generally improve with more trajectories. These findings underscore that both the quality and quantity of offline data, coupled with the GNN-DT architecture, are key to achieving superior performance.
Table 1: Comparison of maximum episode rewards (×$\times 10^{5}$) for baselines and GNN-DT across datasets and context lengths ($K$). Bold indicates the highest value within each dataset and $K$ category.
{tblr}
roweven = c, row3 = c, row5 = c, row7 = c, row9 = c, row11 = c, cell11 = r=2c, cell12 = r=2, cell13 = r=11, cell14 = c=3c, cell17 = c=3c, cell32 = r=3, cell62 = r=3, cell92 = r=3, hline1,12 = -0.08em, hline2 = 4-90.03em, hline3 = 1-2,4-90.03em, hline6,9 = 1-2,4-9dashed, Dataset & Avg. Training
Dataset Reward K=2 K=10
DT Q-DT GNN-DT DT Q-DT GNN-DT
Random $100$ $-2.37$$0.39$ $-1.91$ $-1.97$ $\mathbf{-0.82}$ $-2.12$ $-2.09$ $-1.16$
Random $1000$ $-1.93$ $-2.04$ $-0.86$ $-2.11$ $-2.01$ $-1.18$
Random $10000$ $-1.76$ $-2.04$ $-1.25$ $-1.81$ $-1.98$ $\mathbf{-0.98}$
BaU $100$ $-0.67$$0.07$ $-0.79$ $-0.74$ $\mathbf{-0.59}$ $-0.79$ $-0.72$ $-0.56$
BaU $1000$ $-0.71$ $-0.66$ $-0.65$ $-0.64$ $-0.71$ $-0.57$
BaU $10000$ $-0.69$ $-0.66$ $-0.66$ $\mathbf{-0.44}$ $-0.74$ $-0.53$
Optimal $100$ $-0.01$$0.01$ $-0.67$ $-0.91$ $-0.15$ $-1.12$ $-0.90$ $-0.14$
Optimal $1000$ $-0.63$ $-0.67$ $-0.10$ $-0.87$ $-0.86$ $-0.09$
Optimal $10000$ $-0.63$ $-0.80$ $\mathbf{-0.04}$ $-0.72$ $-0.90$ $\mathbf{-0.07}$
Enhancing Training Datasets
The previous section highlighted that the quality of trajectories in the training dataset is the most influential factor for achieving high performance. In this section, we explore whether creating new datasets by mixing existing ones can further improve performance. The Optimal and Random datasets are combined in different proportions, as summarized in Table 2. A noteworthy result is that supplementing the Optimal dataset with theoretically less useful (Random) trajectories consistently boosts performance. In particular, GNN-DT with $K=10$, trained on a mix of 250 Optimal and 750 Random trajectories, achieves near-oracle results, deviating by only $-0.001\times 10^{5}$ from the optimal reward. A similar trend emerges when blending BaU and Random datasets (Table 8 in Appendix B.4), although the performance gains are not as significant. Overall, these findings indicate that carefully integrating high- and lower-quality data can enhance policy learning beyond what purely Optimal or purely Random datasets can provide.
Table 2: Maximum reward of GNN-DT trained on merged Optimal and Random datasets for $K=2$ and $K=10$. Highest rewards per $K$ are highlighted with bold.
Dataset Total Traj. Avg. Dataset Reward GNN-DT Reward (×$\times 10^{5}$)
K=2 K=10
Random (Rnd.) 100% $1000$ $-2.37$$0.39$ $-0.863$ $-1.187$
Opt. 25% + Rnd. 75% $1000$ $-1.78$$1.07$ $-0.045$ $\mathbf{-0.020}$
Opt. 50% + Rnd. 50% $1000$ $-1.18$$1.19$ $\mathbf{-0.021}$ $-0.040$
Opt. 75% + Rnd. 25% $1000$ $-0.60~{}$±$1.03$ $-0.073$ $-0.057$
Optimal (Opt.) 100% $1000$ $-0.01$$0.01$ $-0.108$ $-0.099$
Component Ablation Study
To better understand the contribution of each architectural component, we conduct an ablation study that systematically removes or replaces elements of our model and then trains the model on the Optimal and Mixed (Opt.25%+Rand.75%). The results in Table 3 reveal that neither a plain DT nor a DT augmented solely with a state-GNN submodule achieves competitive performance. Notably, adding the residual connection atop the state-GNN leads to a significant improvement, from $-0.77\times 10^{5}$ to $-0.16\times 10^{5}$ on the Mixed dataset, demonstrating its importance for effective credit assignment over dynamic inputs. Removing action masking or replacing the GCN module with a Graph Attention Network (GAT) [39] similarly degrades performance, indicating that each component provides distinct and complementary benefits. Ultimately, only the full GNN-DT architecture achieves strong performance across both the Mixed and Optimal datasets.
{tblr}
cells = c, hline1,8 = -0.08em, hline2 = -0.05em, DT & State GNN Action GNN Res. Con. Action Mask Optimal (×$\times 10^{5}$) Mixed (×$\times 10^{5}$)
✓ ✗ ✗ ✗ ✗ $-0.69\pm 0.03$ $-0.95\pm 0.39$
✓ GCN ✗ ✗ ✗ $-0.71\pm 0.02$ $-0.77\pm 0.17$
✓ GCN ✗ ✓ ✗ $-0.18\pm 0.03$ $-0.16\pm 0.07$
✓ GCN GCN ✓ ✗ $-0.11\pm 0.03$ $-0.12\pm 0.04$
✓ GAT GAT ✓ ✓ $-0.14\pm 0.07$ $-0.15\pm 0.06$
✓ GCN GCN ✓ ✓ $\mathbf{-0.09\pm 0.02}$ $\mathbf{-0.10\pm 0.04}$
5.2 Illustrative Example of EV Charging
With the models trained, we proceed to compare the behavior of the best baseline models trained (DT, Q-DT, GNN-DT) against the heuristic BaU algorithm in an EV charging scenario. Fig. 3a presents the SoC progress for three EVs connected one after the other to a single charger throughout the simulation, while Fig. 3b shows the actions of all chargers taken by each algorithm.
(a) EVs’ battery level progression over time.
(b) Charger Actions
In Fig. 3a, the heuristic BaU algorithm consistently overcharges the EVs, often exceeding the desired SoC levels. In contrast, both DT and Q-DT fail to satisfy the desired SoC. Conversely, GNN-DT successfully achieves the desired SoC for all EVs, closely mirroring the behavior of the optimal algorithm. This demonstrates GNN-DT’s ability to precisely control charging based on dynamic state information. Fig. 3b provides further insights into the actions taken by each algorithm. The optimal solution primarily employs maximum charging or discharging power, since it knows the future. In comparison, GNN-DT exhibits a more refined approach, modulating charging power within a range of -6 to 11 kW. Baseline DT and Q-DT display a narrower range of actions, limiting their ability to optimize the charging schedules and adapt to varying conditions. These results underscore the superior capability of GNN-DT in managing the complexities of EV charging dynamics. For a more detailed analysis of key performance metrics in EV charging scenarios, refer to Appendix A.4.
5.3 Evaluation of Generalization and Scalability
Generalization Analysis
(a) Unseen state transition probabilities.
(b) Evaluating on different size problems.
Evaluating the generalization of RL models across varying state transition probabilities is crucial for ensuring consistent performance under diverse conditions [40]. In Fig. 4a, the generalization capabilities of GNN-DT and other baselines are assessed in environments with small, medium, and extreme variations in state transition probabilities (compared to the training environment, see Appendix A.5 for more details about the different distributions). While the baseline methods experience significant performance drops as the evaluation environment deviates from the training setting, GNN-DT maintains strong performance across all scenarios. This highlights the critical role of GNN-based embeddings in improving model robustness and generalization. A key advantage of the GNN-DT architecture, not present in classic DTs, is its invariance to problem size, i.e., the same RL agent can be applied to both smaller and larger-scale environments. Fig. 4b illustrates the scalability and generalization performance of GNN-DT compared to the BaU algorithm and Optimal policy. GNN-DT, trained on a 25-charger setup, was tested on 5, 50, 75, and 100-charger environments. While its performance predictably declines at larger scales, since it wasn’t trained on them, it still outperforms the BaU heuristic, demonstrating robustness to problem-size variation. Training GNN-DT on a mix of charger numbers could potentially further improve its adaptability.
Scalability Analysis
The scalability and effectiveness of GNN-DT were tested when trained on a significantly larger optimization problem involving 250 charging stations. In this scenario, the model must handle up to 250 action variables per step and over 1,000 state variables, which include critical information such as power limits and battery levels. The results presented in Table 4 demonstrate that GNN-DT shows promise for addressing more complex optimization tasks. However, the model requires a substantial increase in both the number of training trajectories and memory resources to maintain efficiency, highlighting a well-known limitation of DT-based approaches. Scaling the problem 10× roughly multiplies GPU memory usage, e.g. storing 3,000 trajectories takes $\approx 2$ GB for 25 chargers versus $\approx 20$ GB for 250. While this can bottleneck large-scale training, parallelization and mini-batching mitigate it, and overall compute scales with the transformer’s context length K (see Fig. 8), pointing to interesting directions for very large problem graphs as future work.
Total Trajectories
Avg. Dataset Reward
GNN-DT Reward
Random $3000$ $-22.39$$1.49$ $-9.34$
BaU $3000$ $-6.67$$0.32$ $-4.23$
Optimal $3000$ $-0.08$$0.03$ $\mathbf{-0.27}$
6 Conclusions
Appendix A Appendix: EV Charging Model

### v3

@@ 1 Introduction
Sequential decision-making problems are critical for efficiently operating a wide array of industries, such as power systems control Roald et al. (2023), logistics optimization Konstantakopoulos et al. (2022), portfolio management Gunjan and Bhattacharyya (2023), and advanced manufacturing processes Gupta and Gupta (2020). However, many practical problems, such as the electric vehicle (EV) charging optimization Panda and Tindemans (2024), are large-scale, have temporal dependencies, and aggregated constraints, often making conventional methods impractical Bubeck (2015). This is especially observed in dynamic environments, where the optimization landscape continuously evolves, requiring real-time solutions.
@@
Reinforcement learning (RL) Sutton and Barto (2018) has been extensively studied for solving optimization problems due to its ability to manage uncertainty, adapt to dynamic environments, and enhance decision-making through trial-and-error Lan et al. (2023); Zhang et al. (2023a). In complex and large-scale scenarios, RL can provide high-quality solutions in real-time compared to traditional mathematical programming techniques that fail to do so Jaimungal (2022). However, RL approaches face significant challenges, such as sparse reward signals that slow learning and hinder convergence to optimal policies Dulac-Arnold et al. (2021). In addition, RL solutions struggle to generalize when deployed in environments different from the one they were trained in, limiting their applicability in real-world scenarios with constantly changing conditions Wang (2024).
@@
Decision Transformers (DT) Chen et al. (2021) is an offline RL algorithm that reframes traditional RL problems as generative sequence modeling tasks conditioned on future rewards Zhang et al. (2023c). DTs utilize previously collected trajectories to inform decision-making without requiring an environment simulator. By learning from historical data, DTs effectively address the sparse reward issue inherent in online RL, relying on demonstrated successful outcomes instead of extensive trial-and-error exploration. However, the trajectory-stitching mechanism of DT often proves insufficient in dynamic real-world environments, leading to suboptimal policies. Although improved variants such as Q-regularized DT (Q-DT) Hu et al. (2024) incorporate additional constraints for greater robustness, they still face significant challenges in generalizing across non-stationary tasks Paster et al. (2024). Consequently, further architectural advances and training strategies are essential to ensure consistent performance in complex environments.
@@
This study introduces GNN-DT [^1] The code can be found at https://github.com/StavrosOrf/DT4EVs., a novel DT architecture that leverages the permutation-equivariant properties of Graph Neural Networks (GNNs) to handle dynamically changing state-action spaces (i.e., varying numbers of nodes over time) and improve generalization. By generating embeddings that remain consistent under node reordering, GNNs offer a powerful way to capture relational information in complex dynamic environments. Moreover, GNN-DT features a novel residual connection between input and output tokens, ensuring that action outputs are informed by the dynamically learned state embeddings for more robust decision-making. To demonstrate the superior performance of the proposed method, we conduct extensive experiments on the complex multi-objective EV charging optimization problem Orfanoudakis et al. (2025a), which encompasses sparse rewards, temporal dependencies, and aggregated constraints. The main contributions are summarized as follows:
@@ 2 Related Work
2.1 Advancements in Decision Transformers
@@
Classic DT encounters significant challenges, including limited trajectory stitching capabilities and difficulties in adapting to online environments. To address these issues, several enhancements have been proposed. The Q-DT Hu et al. (2024) improves the ability to derive optimal policies from sub-optimal trajectories by relabeling return-to-go values in the training data. Elastic DT Wu et al. (2024) enhances classic DT by enabling trajectory stitching during action inference at test time, while Multi-Game DT Lee et al. (2024) advances its task generalization capabilities. The Online DT Zheng et al. (2022); Villarrubia-Martin et al. (2023) extends DTs to online settings by combining offline pretraining with online fine-tuning, facilitating continuous policy updates in dynamic environments. Additionally, adaptations for offline safe RL incorporate cost tokens alongside rewards Liu et al. (2023); Hong et al. (2024). DT has also been effectively applied to real-world domains, such as healthcare Zhang et al. (2023c) and chip design Lai et al. (2023), showcasing its practical utility.
2.2 RL for EV Smart Charging
@@
RL algorithms offer notable advantages for EV dispatch, including the ability to handle nonlinear models, robustly quantify uncertainty, and deliver faster computation than traditional mathematical programming Qiu et al. (2023). Popular methods, such as Deep Deterministic Policy Gradient (DDPG) Jin et al. (2022), Soft Actor Critic (SAC) Jin and Xu (2021), and batch RL Sadeghianpourhamami et al. (2020), show promise but often lack formal constraint satisfaction guarantees and struggle to scale with high-dimensional state-action spaces Yılmaz et al. (2024); Li et al. (2022). Safe RL frameworks address these drawbacks by imposing constraints via constrained MDPs, but typically sacrifice performance and scalability Zhang et al. (2023b); Chen and Shi (2022). Multiagent RL techniques distribute complexity across multiple agents, e.g., charging points, stations, or aggregators Kamrani et al. (2025), yet still encounter convergence challenges and may underperform in large-scale applications. To the best of our knowledge, no study has used DTs to solve the complex EV charging problem, despite DTs’ potential to handle sparse rewards effectively.
3 Problem Formulation